# How digital will the future be? Analysis of prospective scenarios

### v1

@@ Abstract
With the climate change context, many prospective studies, generally encompassing all areas of society, imagine possible futures to expand the range of options. The role of digital technologies within these possible futures is rarely specifically targeted. Which digital technologies and methodologies do these studies envision in a world that has mitigated and adapted to climate change? In this paper, we propose a typology for scenarios to survey digital technologies and their applications in 14 prospective studies and their corresponding 35 future scenarios. Our finding is that all the scenarios consider digital technology to be present in the future. We observe that only a few of them question our relationship with digital technology and all aspects related to its materiality, and none of the general studies envision breakthroughs concerning technologies used today. Our result demonstrates the lack of a systemic view of information and communication technologies. We therefore argue for new prospective studies to envision the future of ICT.
1   Introduction
@@
In the context of climate change, many prospective studies have been proposed in the past years to envision possible futures. They describe the future in very different formats: some are purely narratives, while others rely on quantitative models. These studies help cope with uncertainties and challenges by expanding the available ideas and options  [^1] https://www.geoffmulgan.com/post/social-imagination. Most prospective studies present multiple scenarios (see Figure 1) resulting from different societal or technical hypotheses, each shaped by many variables, such as economic growth, well-being, sobriety, low-carbon energy, population, or uncertain events.
@@
Figure 1: Each study may include several scenarios, representing alternative futures
@@
Prospective scenarios are neither predictions nor forecasts but overviews of possible future paths [Arup, 2019]. Studies often include a baseline business-as-usual scenario used as a reference.
@@
As computer scientists, we wanted to question the place of digital technologies in these studies. Through a preliminary analysis of several scenarios, we observed that they always consider information and communication technologies (ICT) to be present in the future. However, the evolution of the ICT sector and its role in other sectors vary between scenarios. In this paper, we propose to dig further into a set of prospective studies to summarize the future of ICT they envision. More specifically, the objective of this work is to answer the following questions:
@@
  • What ICT is present in the scenarios? i.e., which infrastructure, amount of data, and technologies?
@@
  • How do the scenarios envision the role of ICT for the environmental transition?
To answer these questions, we analyzed 14 studies and their corresponding 35 scenarios, and we focused on the evolution of digital technologies and their areas of applications. We do not pretend to make an exhaustive review of existing scenarios. However, we tried to cover various scenario types, from narrations only to mainly quantitative ones, with positive or negative visions. We identified a limitation of our study, namely the selection of OECD countries-centered studies, which we discuss in Section 5.
@@
Section 2 presents existing work on ICT in future scenarios. Section 3, explains how we selected the studies, and discusses the variables we used to analyze the scenarios. The chosen variables are those that may influence the role of ICT in scenarios. Section 4 more specifically presents what digital technologies look like in the prospective scenarios and their area of applications. Finally, in section 5, we discuss the observations made and identify what we believe are omitted features in the studies.
2   Related works
@@
Only very few works exist regarding the place of digital technologies in prospective scenarios. In this section, we describe the works that are related to ours.
@@
Several authors estimate future ICT energy consumption and carbon footprint [Andrae, 2019, Malmodin and Lundén, 2018]. The projections depend on two main variables: ICT equipment and services efficiency, and demand for ICT (via possible rebound effects) [Freitag et al., 2021]. Our goal here is not to quantify the future consumption of ICT but rather to analyze the possible places of ICT in the future and to explicitly state the implications it could have on its direct and indirect impacts[Hilty et al., 2006] of the sector.
@@
In [Erdmann and Hilty, 2010], the authors propose a scenario analysis of the future impacts of ICT applications on GHG emissions. ICT being a driver of new lifestyles, structural changes, and economic growth, studying its effects is complex. The authors start by reviewing existing macroeconomics on ICT and GHG emissions and conclude that, apart from a few studies, previous works show unambiguous overall reductions of GHG emissions due to ICT. The second part focuses on building specific macroeconomic ICT impact assessment models by looking at ICT’s first-, second-, and third-order effects on the environment. It extends a previous study [Erdmann et al., 2004] by focusing on data uncertainty, requiring further research, and future uncertainty indicating a need for political action. The proposed macroeconomic ICT impact assessment model is applied to the three scenarios from [Erdmann et al., 2004]. These scenarios designed in 2003 describe alternative future courses of ICT until 2020. The scenario analysis from  [Erdmann and Hilty, 2010] concludes with the necessity of incorporating ICT in a general macroeconomic model to quantify rebound effects of energy efficiency measures and the need for a large geographical scope for a complete impact assessment of future scenarios.
@@
Also focusing on assessing future scenarios, authors in [Fauré et al., 2017] review methods for assessing future scenarios’ environmental and social sustainability impacts. They consider three types of scenarios according to the classification of [Börjeson et al., 2006], i.e., predictive, explorative, and normative scenarios. A particular focus is put on sustainability assessments of scenarios in the ICT sector, an example of a rapidly developing sector. The study concludes that the topic of assessment requires further attention if scenarios are to be used to support policy and guide action towards sustainability. In this paper, we do not question the environmental and societal impacts assessment of the scenarios but rather compare scenarios based on the imaginaries and research questions they may drive.
@@
[Andok, 2022] proposes an analysis of the evolution of digital futures in European Union reports. The author reviews three EU reports on digital futures. The first one, [European Commission, 2014], published in 2014, is very techno-optimistic in every aspect of the society. The second (post-COVID pandemic) report [European Commission, 2021] from 2021 is slightly less optimistic and mentions vulnerabilities and disinformation. It also discusses sustainability, mentioning how digital technologies can contribute to EU Green Deal objectives. The third report [Polona and Stefano, 2022] gives short-term objectives regarding cyber-security after the beginning of the Russian-Ukrainian war. From this analysis, Andok highlights the shift from opportunities only to dangers and threats, and from 40 years perspective visions to shorter ones (10 years or even less).
3   Methodology
This section describes the 14 studies (section 3.2) we chose to analyze. We also present the features we use to classify them, which can influence on the evolution of ICT within the different scenarios (section 3.1).
3.1 Typologies describing the studies and scenarios
@@
Each study has its way of designing possible futures, so capturing the diversity of the scenarios is complex [Van Notten et al., 2003]. Several works have proposed scenario typologies (e.g. [Ducot and Lubben, 1980, Van Notten et al., 2003, Crawford, 2019]) to capture this diversity. These typologies aim to categorize scenarios by answering the questions of why (project goal), how (process design), and what (scenario content) [Van Notten et al., 2003]. However, more analysis is needed to understand more deeply why scenarios end up being so similar or different. Scenarios result from hypotheses and parameters evolutions regarding economy, lifestyles and governance, which are not always explicit.
@@
In this work, we focus on the place and role of ICT in future perspectives. We are therefore interested in comparing the different scenarios according to features that influence the place of ICT. Authors [Erdmann et al., 2004, Erdmann and Hilty, 2010] distinguish external variables that include GDP, population, labor demand, electricity prices, and other variables that may influence first-, second-, and third-order impacts of ICT. Examples of variables impacting first-order impacts are technological trends from large to small devices, household internet penetration, or energy efficiency. Variables for second-order impacts include energy efficiency or materials potentials for other sectors of activity. Finally, variables for third-order impacts are elasticity of demand, average personal transport time, or some economic behavior. The authors use all these variables to analyze the three scenarios from [Erdmann et al., 2004].
@@
Typology
Level:
Study
Objective:
General overview of the study
Features overview:
Project goal, content and complexity
Reference:
[Van Notten et al., 2003]
Societal variables
Level:
Scenario
Objective:
non ICT-specific variables that influence scenarios
Features overview:
Economy, lifestyle, demography, governance
ICT variables
Level
Scenario
Objective
ICT-specific variables present in scenarios
Features overview
Infrastructure, Applications, Technologies
@@
Table 1: Overview of typology and parametric variables used to compare scenarios
@@
The first one is the general typology from [Van Notten et al., 2003]. It enables to view how the study was conducted and its perimeter quickly. The second one is Societal variables, which concern hypotheses for society pathways regarding lifestyle, economy, governance, demography, and equality. We propose variables freely adapted from  [Ademe, 2021]. When analyzing the scenarios, we observed that the variables of economic growth, sobriety, the importance of the state, and the collaboration between countries were discussed in all general studies and drive scenarios. We here refer to general studies as the ones that regard all areas of society. These societal variables impact the kind of ICT envisioned; for example, a scenario with solid states implies centralization of digital infrastructure and services.
@@
Finally, we considered ICT-specific variables to characterize the ICT infrastructure and usage types. We distinguished three types of ICT-specific variables, the first being general ICT characteristics, with variables about the evolution of usages, about equipment (lifetimes, quantity of equipment, data center locations, use of cloud,…), and about usage (energy efficiency, data flows and data center consumption). All these characteristics directly impact equipment and infrastructures, and we call them infrastructure variables. The second type, called applications, addresses the kind of applications of ICT considered in the scenarios (teleworking, smart mobility, robots for manufacturing in industry,…), and the third type, technologies, the digital technologies present in the scenario (data analytics, AI, cloud and edge computing,…).
@@
Even if digital technologies exist in all scenarios, the studies rarely mention these variables. We observed that they are only discussed explicitly in quantitative studies that calculate energy consumption and greenhouse gas emissions (e.g., [Ademe, 2021, négaWatt, 2021, RTE, 2022]) and in prospective studies specifically targetting ICT as main topic (e.g., [Creutzig et al., 2022, CNIL, 2021, Deron and McDonald, 2022]). In fact, as previously observed by [Creutzig et al., 2022], no current conceptualization of decarbonization pathways explicitly accounts for the impacts of digitalization in the Anthropocene.
Figure 3 gives the complete list of variables.
3.2 Studies selection
In this section, we present how we chose the 14 studies and how we categorized them according to a subset of features that we found to be relevant. We give a short overview of the considered studies in Appendix A. We also provide a complete classification of all scenarios according to the typology and variables described in the previous section in the supplementary material.
We first selected studies we already knew, then searched for others online. We tried to find studies with various: spatial perimeters (which countries are considered), domains considered (ICT -specific or not), and methodologies (narratives vs quantitative studies). Other variables, such as those of the typology (section 3.1), also help differentiate scenarios but were not our main starting points. The last criterion we used, which is not part of the typology from [Van Notten et al., 2003], is the publication year.
Publication year
We limited ourselves to recent studies. The oldest one in our deep analysis was published in 2018. We found very few notions of digital technologies in the oldest general scenarios. For example, in  [OECD, 2011], we found only one mention of data access, smart metering, and a suggestion to make chemical information widely available through the Internet thanks to the development of new and innovative computer models. Another example of a scenario that we purposely did not include in this work is the ICT-specific from [Erdmann et al., 2004]. It aims to explore (qualitatively) and assess (quantitatively) how ICT will influence environmental sustainability between 2003 and 2020.
Spatial scale
As we are located in France, we naturally started our review with the most recent prospective works made in this country: [Ademe, 2021, négaWatt, 2021, Ministère de la Transition écologique et solidaire, 2020, RTE, 2022, The Shift Project, 2020, Millot et al., 2018, CNIL, 2021]. We then opened to studies targeting the future of EU: [European Commission, 2019, European Environment Agency, 2021]. Finally, we selected studies for the entire world: [IPCC, 2022, Arup, 2019, Danish Design Center, 2020, Creutzig et al., 2022]. One ICT-specific study focuses on Quebec [Deron and McDonald, 2022]. The studies chosen thus mostly target OECD countries, which is a limitation of our work.
Data
With this criterion, we differentiate qualitative only from more quantitative studies. Within qualitative, we include some entirely narrative studies: [European Environment Agency, 2021, Arup, 2019, Danish Design Center, 2020]. They include stories of people living in 2050 enabling us to look at what life could more concretely look like, and resemble science fiction. The study [Danish Design Center, 2020] is in the form of a future scenario toolkit. Many other toolkits [^2] https://oecd-opsi.org/toolkits/?_toolkit_discipline_or_practice=futures-and-foresight exist but were not studied here as none are ICT-centered, and we considered these toolkits to be redundant with the ones already studied. Conversely, some studies are much more quantitative and try to estimate GHG emissions by 2050 (e.g., [RTE, 2022] is mainly based on numerical models).
Topic
All studies build upon climate change context. General prospective studies concern all societal areas (e.g. [Ademe, 2021, IPCC, 2022, The Shift Project, 2020, Arup, 2019]. Additionally, we selected two studies focusing on energy in France in 2050 [RTE, 2022, Millot et al., 2018]. Finally, obviously, we have included in our work studies that specifically target information and communication technologies: [Creutzig et al., 2022, CNIL, 2021, Deron and McDonald, 2022].
4   Results

### v3

@@ Abstract.
Within the context of climate change, many prospective studies generally encompassing all areas of society, imagine possible futures to expand the range of options. The role of digital technologies within these possible futures is rarely specifically targeted. Which digital technologies and methodologies do these studies envision in a world that has mitigated and adapted to climate change? In this paper, we propose a typology for scenarios to survey digital technologies and their applications in 14 prospective studies and their corresponding 35 future scenarios. Our finding is that all the scenarios consider digital technology to be present in the future. We observe that only a few of them question our relationship with digital technology and all aspects related to its materiality, and none of the general studies envision breakthroughs concerning technologies used today. Our result demonstrates the lack of a systemic view of information and communication technologies. We therefore argue for new prospective studies to envision the future of ICT.
future of ICT, prospective studies, future scenarios, sustainable ICT
copyright: noneconference: Tenth Workshop on Computing within Limits; June 18–19, 2024; Online
1. Introduction
@@
In the context of climate change, many prospective studies have been proposed in recent years to envision possible futures via both narrative and quantitative formats. Different techniques may be used in future study: reflection, simulation and modeling, monitoring or behavioral (Mankoff et al., 2013). Prospective studies often present multiple scenarios (see Figure 1) resulting from different societal or technical hypotheses, each shaped by many variables, such as economic growth, well-being, sobriety, low-carbon energy, population, or uncertain events. They help cope with uncertainties and challenges by expanding the horizons of futures that can be imagined  [^1] https://www.geoffmulgan.com/post/social-imagination. Future studies can help being better prepared for meeting future challenges, enabling the invention of a better future, or increasing the long-term impact of our work (Mankoff et al., 2013). They are actively used by policy designers.
@@
Figure 1. Each study may include several scenarios, representing alternative futures
@@
The prospective scenarios are neither predictions nor forecasts but rather overviews of possible future paths (Arup, 2019). Studies often include a baseline business-as-usual scenario used as a reference.
The LIMITS community has a long history of working on future studies and on ICT futures (Mankoff et al., 2013; Penzenstadler et al., 2014; Nardi, 2016; Tanenbaum et al., 2016; Eriksson and Pargman, 2018; Pargman et al., 2019). Among the many research topics they address, the community questions what will research look like in the context of scarcity, how design fiction can challenge business-as-usual and pluralize the future, how fiction allows to adopt a range of different intellectual commitments and values about the future, and how to ascend from future studies to the concrete. Authors of (Nardi et al., 2018) underline the role that computing has to play in responding to global limits and in shaping a society that meaningfully adapts to them. They nevertheless state that many computing researchers and practitioners in practice assume there is only one possible likely future, one very much like the present.
@@
As computer scientists, we wanted to question the place of digital technologies in studies that have recently been designed to support policy and guide action towards sustainability, and to develop the imagination in this context. Through a preliminary analysis of several scenarios, we observed that they always consider Information and Communication Technologies (ICT) to be present in the future. However, the evolution of the ICT sector and its role in other sectors vary between scenarios. In this paper, we propose to dig further into a set of prospective studies to summarize the future of ICT they envision. More specifically, the objective of this work is to answer the following questions:
@@
  • What ICT is present in the scenarios, i.e., which infrastructure, amount of data, and technologies?
@@
To answer these questions, we analyzed 14 studies and their corresponding 35 scenarios, and we focused on the evolution of digital technologies and their areas of applications. We do not pretend to make an exhaustive review of existing scenarios. However, we tried to cover several scenario types within a spectrum of narrative and quantitative formats and present positive or negative visions of the future. We identified a limitation of our study, namely the selection of OECD countries-centered studies, which we discuss in Section 5.
@@
Section 2 presents existing work on ICT concerning future scenarios. Section 3 explains how the studies were selected and discusses the variables used to analyze the scenarios. Section 4 details the results for ICT variables. A focus is made on the digital technologies and area of applications explicitely mentioned in the scenarios. Lastly, in section 5, we discuss the omitted features in the studies that we identified.
2. Related works
@@
We found only very few studies that compare and analyze the place of digital technologies in prospective scenarios. In this section, we describe the works that we found related to ours.
@@
Several authors estimate future ICT energy consumption and carbon footprint (Andrae, 2019; Malmodin and Lundén, 2018). The projections depend on two main variables: ICT equipment and services efficiency, and demand for ICT (via possible rebound effects) (Freitag et al., 2021). Our goal here is not to quantify the future consumption of ICT but rather to analyze the possible places of ICT in the future and to explicitly state the implications it could have on first-, second-, and third-order effects (Hilty et al., 2006) of the sector.
@@
In (Erdmann and Hilty, 2010), the authors propose a scenario analysis of the future impacts of ICT applications on GHG emissions. ICT being a driver of new lifestyles, structural changes, and economic growth, studying its effects is complex. The authors start by reviewing existing macroeconomics on ICT and GHG emissions and conclude that, apart from a few studies, previous works show unambiguous overall reductions of GHG emissions due to ICT. The second part focuses on building specific macroeconomic ICT impact assessment models by looking at ICT’s first-, second-, and third-order effects on the environment. It extends a previous study (Erdmann et al., 2004) by focusing on data uncertainty, requiring further research, and future uncertainty indicating a need for political action. The proposed macroeconomic ICT impact assessment model is applied to the three scenarios from (Erdmann et al., 2004). These scenarios designed in 2003 describe alternative future courses of ICT until 2020. The scenario analysis from  (Erdmann and Hilty, 2010) concludes with the necessity of incorporating ICT in a general macroeconomic model to quantify rebound effects of energy efficiency measures and the need for a large geographical scope for a complete impact assessment of future scenarios.
@@
Also focusing on assessing future scenarios, authors in (Fauré et al., 2017) review methods for assessing future scenarios’ environmental and social sustainability impacts. They consider three types of scenarios according to the classification of (Börjeson et al., 2006), i.e., predictive, explorative, and normative scenarios. A particular focus is put on sustainability assessments of scenarios in the ICT sector, an example of a rapidly developing sector. The study concludes that the topic of assessment requires further attention if scenarios are to be used to support policy and guide action towards sustainability. In this paper, we do not question the environmental and societal impacts assessment of the scenarios but rather compare scenarios based on the imaginaries and research questions they may drive.
@@
Questioning the evolution in the vision for ICT in European policies, (Andok, 2022) proposes an analysis of digital futures in different European Union reports. The author reviews three EU reports on digital futures. The first one, (European Commission, 2014), published in 2014, is very techno-optimistic in every aspect of the society. The second (post-COVID pandemic) report (European Commission, 2021) from 2021 is slightly less optimistic and mentions vulnerabilities and disinformation. It also discusses sustainability, mentioning how digital technologies can contribute to EU Green Deal objectives. The third report (Polona and Stefano, 2022) gives short-term objectives regarding cyber-security after the beginning of the Russian-Ukrainian war. From this analysis, Andok highlights the shift from opportunities only to dangers and threats, and from 40 years perspective visions to shorter ones (10 years or even less).
Finally, in (Penzenstadler et al., 2014), the authors present and analyze a compilation of fictional abstracts written by researchers attending the ICT4S conference. They cluster the abstracts according to four major aspects (infrastructure, society, food, and waste), before stating major changes suggested in technological, economic, and social dimensions. This work is related to ours as it aimed at finding links between different scenarios.
3. Methodology
This section describes the 14 studies (section 3.1) we chose to review. From these studies and the state of the art on prospective analysis, we identified features that can influence the evolution of ICT within the different scenarios and organized them in a typology as described in section 3.2. The 14 studies were then classified according according to the proposed methodology. While we briefly provide first results of the analysis in this section, next section provides a detailed analysis on the representation of ICT the scenarios.
3.1. Studies selection
In this section, we present how we chose the 14 studies. We give a short overview of the considered studies in Appendix A.
We first selected studies we already knew, then searched for others online. We tried to find studies with various: spatial perimeters (which countries are considered), domains considered (ICT -specific or not), and methodologies (narratives vs quantitative studies). The last criterion we used is the publication year.
Topic
All studies build upon climate change context. We refer to general studies as the ones that regard all areas of society. We started our analysis with such general studies (e.g. (Ademe, 2021; IPCC, 2022; Shift Project, 2020; Arup, 2019)). In this work, nine studies are general. Additionally, we selected two studies focusing on energy in France in 2050 (RTE, 2022; Millot et al., 2018). Finally, obviously, we have included in our work studies that specifically target information and communication technologies: (Creutzig et al., 2022; CNIL, 2021; Deron and McDonald, 2022). Although we did not select the studies according to the presence of ICT in them, all of them actually mention ICT.
Publication year
We limited ourselves to recent studies. The oldest one in our deep analysis was published in 2018. We found very few notions of digital technologies in the oldest general scenarios. For example, in  (OECD, 2011), we found only one mention of data access, smart metering, and a suggestion to make chemical information widely available through the Internet thanks to the development of new and innovative computer models. Another example of a scenario that we purposely did not include in this work is the ICT-specific from (Erdmann et al., 2004). It aims to explore (qualitatively) and assess (quantitatively) how ICT will influence environmental sustainability between 2003 and 2020.
Spatial scale
As we are located in France, we naturally started our review with the most recent prospective works made in this country: (Ademe, 2021; négaWatt, 2021; Ministère de la Transition écologique et solidaire, 2020; RTE, 2022; Shift Project, 2020; Millot et al., 2018; CNIL, 2021). We then opened to studies targeting the future of EU: (European Commission, 2019; European Environment Agency, 2021). Finally, we selected studies for the entire world: (IPCC, 2022; Arup, 2019; Danish Design Center, 2020; Creutzig et al., 2022). One ICT-specific study focuses on Quebec (Deron and McDonald, 2022). The studies chosen thus mostly target OECD countries, which is a limitation of our work.
Data
With this criterion, we differentiate qualitative only from more quantitative studies. Within qualitative, we include some entirely narrative studies: (European Environment Agency, 2021; Arup, 2019; Danish Design Center, 2020). They include stories of people living in 2050 enabling us to look at what life could more concretely look like, and resemble science fiction. The study (Danish Design Center, 2020) is in the form of a future scenario toolkit. Many other toolkits [^2] https://oecd-opsi.org/toolkits/?_toolkit_discipline_or_practice=futures-and-foresight exist but were not studied here as none are ICT-centered, and we considered these toolkits to be redundant with the ones already studied. Conversely, some studies are much more quantitative and try to estimate GHG emissions by 2050 (e.g., (RTE, 2022) is mainly based on numerical models).
3.2. Typologies describing the studies and scenarios
@@
Each study has its way of designing possible futures, so capturing the diversity of the scenarios is complex (Van Notten et al., 2003). Several works have proposed scenario typologies (e.g. (Ducot and Lubben, 1980; Van Notten et al., 2003; Crawford, 2019)) to capture this diversity. These typologies aim to categorize scenarios by answering the questions of why (project goal), how (process design), and what (scenario content) (Van Notten et al., 2003). However, more analysis is needed to understand more deeply why scenarios end up being so similar or different. Scenarios result from hypotheses and parameters evolutions regarding economy, lifestyles and governance, which are not always explicit.
In this work, we focus on the place and role of ICT in future perspectives. We are therefore interested in comparing the different scenarios from our selected studies according to features that influence the place of ICT.
@@
Authors (Erdmann et al., 2004; Erdmann and Hilty, 2010) distinguish external variables that include GDP, population, labor demand, electricity prices, and other variables that may influence first-, second-, and third-order impacts of ICT. Examples of variables impacting first-order impacts are technological trends from large to small devices, household internet penetration, or energy efficiency. Variables for second-order impacts include energy efficiency or materials potentials for other sectors of activity. Finally, variables for third-order impacts are elasticity of demand, average personal transport time, or some economic behavior. The authors use all these variables to analyze the three scenarios from (Erdmann et al., 2004). These variables focus on impacts of ICT but do not give a clear picture of what ICT is present in the scenarios and the place of ICT if the futures.
@@
Goal and design Level: Study
Objective: General overview of the study
Features overview: Project goal, content and complexity
Reference: (Van Notten et al., 2003)
Societal variables Level: Scenario
Objective: non ICT-specific variables that influence scenarios
Features overview: Economy, lifestyle, demography, governance
ICT variables Level Scenario
Objective ICT-specific variables present in scenarios
Features overview Infrastructure, Applications, Technologies
@@
Table 1. Overview of typology and parametric variables used to compare scenarios
@@
The first one is the general typology from (Van Notten et al., 2003), that we call ”goal and design”. It enables to view how the study was conducted and its perimeter. The second one is ”Societal variables”, which concern hypotheses for society pathways regarding lifestyle, economy, governance, demography, and equality. We propose variables freely adapted from  (Ademe, 2021). When analyzing the scenarios, we observed that the variables of economic growth, sobriety, the importance of the state, and the collaboration between countries were discussed in all general studies and drive scenarios. These societal variables impact the kind of ICT envisioned; for example, a scenario with solid states may favor centralization of digital infrastructure and services.
@@
Finally, after reading the different scenarios we selected and organized ICT-specific variables to characterize the ICT infrastructure and usage types. We distinguished three types of ICT-specific variables, the first being general ICT characteristics, with variables about the evolution of usages, about equipment (lifetimes, quantity of equipment, data center locations, use of cloud,…), and about usage (energy efficiency, data flows and data center consumption). All these characteristics directly impact equipment and infrastructures, and we call them infrastructure variables. The second type, called applications, addresses the kind of applications of ICT considered in the scenarios (teleworking, smart mobility, robots for manufacturing in industry,…), and the third type, technologies, the digital technologies present in the scenario (data analytics, AI, cloud and edge computing,…).
@@
Even if digital technologies exist in all scenarios, the studies rarely mention ICT-specific variables as driving the scenarios pathway. We observed that they are only discussed explicitly in quantitative studies that calculate energy consumption and greenhouse gas emissions (e.g., (Ademe, 2021; négaWatt, 2021; RTE, 2022)) and in prospective studies specifically targetting ICT as main topic (e.g., (Creutzig et al., 2022; CNIL, 2021; Deron and McDonald, 2022)). In fact, as previously observed by (Creutzig et al., 2022), no current conceptualization of decarbonization pathways explicitly accounts for the impacts of digitalization in the Anthropocene.
Figure 3 gives the complete list of variables. The encoding of all scenarios according to our typology and variables is provided in the supplementary material  [^3] https://hal.science/hal-04362220v2/file/Supplementary_Analysis%20of%20ICT%20in%20prospective%20scenarios.pdf.
4. Results for ICT variables